OpenStudy (anonymous):

Trini's car breaks down on a highway. Trini estimates that she is 20 to 30 miles from the nearest car repair shop. She calls a towing company that charges a fee of $80 plus $3 per mile to tow a car. If Trini uses this towing company, which is the best estimate for the amount of money, m, she will pay for the company to tow her car?

OpenStudy (anonymous):

$80 + 3*30 = $80+$90 = $170 (maximum)

OpenStudy (anonymous):

Yep, m = 80 + 3 * 30 = $170 charge for towing up to 30 miles.